Jess pays $2510 per quarter over 3 years on a flat rate loan of $24000.
a)Calculate the total interest paid.
b)Calculate the annual interest rate.

Answer by Theo(13342)

flat rate loan is 24000.
jesse pays 2510 per quarter over 3 years.
her total payments are 2510 * 4 * 3 = 30120
her interest paid is 30120 - 24000 = 6120
her flat rate interest rate is 6120/3/24000 = .085 = 8.5% per year.

24000 * .085 * 3 = total interest of 6120.
total payment due by the end of the loan is 24000 + 6120 = 30120
this is divided into 12 equal payments of 2510 each quarter.

the flat rate loan is the same as the simple interest type loan.

the formula is i = p * r * n

p is the principal
r is the interest rate per time period
n is the number of time periods.

in your problem.

p = 24000
r = .085 / 4 = .02125 per quarter.
n = 3 * 4 = 12 quarters.

i = p * r * n becomes i = 24000 * .02125 * 12 = 6120

that's the total interest due on a principal of 24000.

add that to the principal and the total amount due by the end of the loan is 24000 + 6120 = 30120.